# Leveraging Large Language Models for Career Mobility Analysis: A Study of Gender, Race, and Job Change Using U.S. Online Resume Profiles

### v1

@@ 10.7 Authors’ contributions
PA conceptualized the research, developed the methodology, implemented the software, conducted the investigation, performed formal analysis and validation, and was responsible for visualization, writing the original draft, and manuscript review and editing. CX and XJSA contributed to software development and data curation. YL assisted in conceptualization, methodology design, validation, and manuscript review and editing. EPL provided conceptualization, methodology refinement, validation, resources, supervision, project administration, funding acquisition, and manuscript review and editing. All authors read and approved the final manuscript.
@@ Appendix A Race Classification Evaluation
We acknowledge the lower prediction accuracy for Black and Asian names. Nevertheless, this measurement error in the inferred race category variable is expected to impact the main regression analysis in a specific and predictable way. Crucially, measurement error in an independent variable (the inferred race) tends to attenuate estimated coefficients toward zero, rather than introduce spurious associations This attenuation bias (also known as regression dilution bias) is well-documented. Furthermore, systematic misclassification in name-based race inference methods has been observed, with studies indicating that such errors are correlated with demographic and socioeconomic factors [26]. Taken together, these findings suggest that any observed race-related effects in our regression analyses are likely conservative estimates. Readers should interpret the magnitude of these coefficients with this caution in mind.
@@ Appendix B Benchmarking Recent LLMs for SOC Classification
Therefore, for this experiment, we utilized the LLM-assisted ground truth SOC labels constructed by Achananuparp et al. [1], leveraging the same Jobs12K dataset. This approach utilizes GPT-4o as an expert annotator (LLM-as-a-Judge) to select the final ground truth from a combined pool of predicted candidates. The final ground truth labels represent the majority vote of three independent GPT-4o runs, where the model was tasked with selecting the most appropriate SOC label(s) from a shuffled list of all candidates generated by the tested methods.
@@
The accuracy scores for all models on the Jobs12K test set are presented in Table 6. Firstly, Gemini 2.5 Flash achieved the highest accuracy (0.7425), outperforming the GPT-3.5 Turbo benchmark by 2.3 percentage points. Despite being a newer efficient GPT-based model, GPT-4.1 Mini (0.6073) performed significantly poorer than GPT-3.5 Turbo (0.7195). This unexpected result suggests that the architectural trade-offs made to optimize GPT-4.1 Mini may have compromised the model’s ability to handle SOC classification. Lastly, The Llama 3.1 8B Instruct model lagged substantially behind all the other models, achieving an accuracy of 0.4722. This finding is consistent with recent studies [1], which similarly showed the comparatively poor performance of Llama models in the SOC classification task.

### v2

@@ 10.7 Authors’ contributions
PA conceptualized the research, developed the methodology, implemented the software, conducted the investigation, performed formal analysis and validation, and was responsible for visualization, writing the original draft, and manuscript review and editing. YX and XJSA contributed to software development and data curation. YL assisted in conceptualization, methodology design, validation, and manuscript review and editing. EPL provided conceptualization, methodology refinement, validation, resources, supervision, project administration, funding acquisition, and manuscript review and editing. All authors read and approved the final manuscript.
@@ Appendix A Race Classification Evaluation
We acknowledge the lower prediction accuracy for Black and Asian names. Nevertheless, this measurement error in the inferred race category variable is expected to impact the main regression analysis in a specific and predictable way. Crucially, measurement error in an independent variable (the inferred race) tends to attenuate estimated coefficients toward zero, rather than introduce spurious associations This attenuation bias (also known as regression dilution bias) is well-documented. Furthermore, systematic misclassification in name-based race inference methods has been observed, with studies indicating that such errors are correlated with demographic and socioeconomic factors [kozlowski2022avoiding]. Taken together, these findings suggest that any observed race-related effects in our regression analyses are likely conservative estimates. Readers should interpret the magnitude of these coefficients with this caution in mind.
@@ Appendix B Benchmarking Recent LLMs for SOC Classification
Therefore, for this experiment, we utilized the LLM-assisted ground truth SOC labels constructed by achananuparp2025multi, leveraging the same Jobs12K dataset. This approach utilizes GPT-4o as an expert annotator (LLM-as-a-Judge) to select the final ground truth from a combined pool of predicted candidates. The final ground truth labels represent the majority vote of three independent GPT-4o runs, where the model was tasked with selecting the most appropriate SOC label(s) from a shuffled list of all candidates generated by the tested methods.
@@
The accuracy scores for all models on the Jobs12K test set are presented in Table 6. Firstly, Gemini 2.5 Flash achieved the highest accuracy (0.7425), outperforming the GPT-3.5 Turbo benchmark by 2.3 percentage points. Despite being a newer efficient GPT-based model, GPT-4.1 Mini (0.6073) performed significantly poorer than GPT-3.5 Turbo (0.7195). This unexpected result suggests that the architectural trade-offs made to optimize GPT-4.1 Mini may have compromised the model’s ability to handle SOC classification. Lastly, The Llama 3.1 8B Instruct model lagged substantially behind all the other models, achieving an accuracy of 0.4722. This finding is consistent with recent studies [achananuparp2025multi], which similarly showed the comparatively poor performance of Llama models in the SOC classification task.